OpenStudy (anonymous):

solve for t. l=980t^2/4π^2

OpenStudy (anonymous):

\[I=\frac{980t^2}{4\pi^2}\] like that?

OpenStudy (anonymous):

multiply both sides by \(4\pi^2\) to get \[4\pi^2I=980t^2\] divide by \(980\) to get \[\frac{\pi^2I}{245}=t^2\] and finally take the square root of both sides, don't forget the \(\pm\)

OpenStudy (anonymous):

so im still not 100% sure what the answer is. could you maybe show me the last step?